OpenStudy (anonymous):

HELP! WILL GIVE MEDAL! Given the two expressions shown below: A: Square root of 4 + square root of 25 B: Square root of 4 + square root of 2 ------------- A: Both are rational B: Both are irrational C: A is rational, but B is not irrational D: A is irrational, but B is rational Is C correct?

OpenStudy (anonymous):

\[A: \sqrt{4}+\sqrt{25} ------ B: \sqrt{4}+\sqrt{2}\]

OpenStudy (saldis):

yes it is 

OpenStudy (saldis):

Correct girl

OpenStudy (mathmate):

@Shiva_ Is \(\sqrt 2\) rational?

OpenStudy (michele_laino):

@Shiva_ please note that: \[4=2^{2},25=5^{2}\]

OpenStudy (saldis):

(i.e. as a ratio).

OpenStudy (saldis):

2 is not rational

OpenStudy (saldis):

dont think sooooo

OpenStudy (michele_laino):

@saldis please note that \[2=\frac{ 2 }{ 1 }\] so 2 is a rational number

OpenStudy (saldis):

sorry yes it is just worked it out :)

OpenStudy (michele_laino):

@saldis you are welcome!:)

OpenStudy (saldis):

thnks your amazing !!!!!! :)